OpenStudy (anonymous):

complete the table 1 x^2 dx Rewrite: Integrate: Simplify:

OpenStudy (aravindg):

\[\large \int\limits x^n=\dfrac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

\[\int\limits_{}^{}x ^{2}dx=x ^{3}/3\]

OpenStudy (zehanz):

\[\int\limits_{}^{}x^2dx=\frac{ 1 }{ 3 }x^3+C\]